Draw the condensed structural formula for the product obtained from the hydrogenation of glyceryl tripalmitoleate.
Answer #1

CH₂-O- & fchal 7 CH=CH-CH2) CH₃ -C-(CH2)₂ - CH=CH-(CH2) 5-CH3 + 3H₂ CH₂-O-C-(CH2)₂ -CH=CH-(CH2) 5 CH3 Glyceryl toipalmitolea